40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9673     Entry time: Tue Feb 25 17:27:41 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: REFL signals calibrated 

I have recalibrated the REFL signals.

I first adjusted the demod phases until the I-signals lined up with the I-phase in the sensing matrix plot:

SensMat_25Feb2014.png

I then balanced the ITM drives by pushing on -1*ITMX and +1.015*ITMY, and seeing a minimum of MICH actuation in the I-phase of REFL55 (the PD I was locking with).

I then took a nice long measurement with DTT, and measured the peak heights in I and Q for each REFL diode.  I was driving PRM with 100 cts at 675.1Hz, and ITMX with 1000 cts at 452.1 Hz (and matching ITMY drive, to make pure MICH).  Knowing these numbers, and the actuator calibrations (PRM elog 8255, ITMs elog 8242), I know that I was driving PRCL by ~4.3 pm, and MICH by ~23 pm. 

For the I-phase calibrations, I find the peak height at the PRCL drive frequency, and divide 4.3 pm by that height.  For the Q-phase calibrations, I find the peak height at the MICH drive frequency, and divide 23 pm by that height.

This gives me the following calibrations:

  Calibration [picometers / count]
REFL 11 I    0.15
REFL 11 Q   21.6
REFL 33 I    1.06
REFL 33 Q  209
REFL 55 I    0.9
REFL 55 Q   27       
REFL 165 I    0.1
REFL 165 Q   11.6

 My calibrated REFL spectra then looks like:

Calibrated_25Feb2014.pdf

ELOG V3.1.3-